OpenStudy (anonymous):

I WILL FAN YOU AND GIVE YOU A MEDAL Part 2: What is the simplified expression? (3x3y4)2

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

none of your business

OpenStudy (anonymous):

joking

OpenStudy (anonymous):

loll

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

@julian25

OpenStudy (jhannybean):

\[\large (3x^3y^4)^2\]distribute the squared into all terms. \[\large (3)^2\cdot (x^3)^2\cdot (y^4)^2\]\[\large 3^2 \cdot x^{3\cdot 2} \cdot y^{4\cdot 2}\]

OpenStudy (jhannybean):

And now just simplify.

OpenStudy (anonymous):

9X^6Y^8

OpenStudy (anonymous):

Thats what i'm getting but i thought they wanted me to do like x^6 xxxxxx

OpenStudy (jhannybean):

Nah. That's your simplified answer :)

OpenStudy (jhannybean):

Good job :)

OpenStudy (anonymous):

Loll thank you very much have a goodnight and let me fan you and give you a medal :D